OpenStudy (anonymous):

what irrational number between 7.7 and 7.9 ? i need to find a irrational number between those 2

OpenStudy (slinley):

i dont knowww. wait hold on.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

7/9 would be irrational, I think.

OpenStudy (anonymous):

hmmmm im not getting this at all

OpenStudy (anonymous):

Well, 7/9 cannot be expressed as a decimal, since it is 7.7777 never ending.

OpenStudy (slinley):

yes, 7/9 but want an explanation??

OpenStudy (anonymous):

yes please!!!!!!!

OpenStudy (anonymous):

wait but isnt 7.77777 repeating a irrational number right

OpenStudy (slinley):

yes, ur right, but lemme give u an explanation!

OpenStudy (anonymous):

ok

OpenStudy (slinley):

irrational means not rational, so ........

OpenStudy (slinley):

hold on, im typing moree

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

7.8 Wouldn't be an irrational number because it can be written as a simple fraction 39/5

OpenStudy (anonymous):

An irrational number is one that doesn't have a simple fraction form.

OpenStudy (anonymous):

well then what is there between 7.7 and 7.9 thats a irrational number cuz 7.8 s the only number between them

OpenStudy (slinley):

yes, ty Ankh...thats exactly what i was going to say :)

OpenStudy (slinley):

good job @bailzbear :)

OpenStudy (anonymous):

sooooooooooooooooooooooooooooooooooo.......... the answer is 7.9 right

OpenStudy (slinley):

ummm. lemme check ;p

OpenStudy (anonymous):

7.9 = 79/10 = rational.

OpenStudy (slinley):

yassss

OpenStudy (anonymous):

\[\sqrt{60} = 7.745966692 = Irrational\]

OpenStudy (anonymous):

YAY I GOT IT!!!!!!!!!! \(^-^)/

OpenStudy (anonymous):

So the square root of 60 would be an appropriate answer.

OpenStudy (anonymous):

thank you so much

OpenStudy (slinley):

can you medall me/ or us :P ??? For helping yhuu or fan us

OpenStudy (slinley):

lmaoooooooooooooooooooooooooooooooooooooooooooooooooooooooo